# Shift orbifolds, decompactification limits, and lattices

### v1

@@ 2.1 Lattice set up
Specializing to $d=1$, we denote the heterotic moduli by a pair $(r,{{A}})$, where ${{A}}\in{\mathbb{R}}^{16}$ is the set of Wilson line parameters. A choice of $(r,{{A}})$ determines a point $\rho$ in the Grassmannian $\operatorname{Gr}(1,17)$ through the orthogonal basis for ${\mathbb{R}}^{1,17}$ consisting of orthogonal vectors ${\widetilde{{\pi}}}$ , ${\pi}$, and ${\pi}^{\circ}_{\alpha}$, $\alpha=1,\ldots,16$
@@ 2.2 Vertex operators
The heterotic worldsheet theory consists of the Narain CFT with $d+16$ left-moving chiral bosons $X^{I}_{\text{\tiny{L}}}(z)$, ${\cal X}_{\text{\tiny{L}}}^{a}(z)$, and $d$ right-moving chiral bosons $X^{I}_{\text{\tiny{R}}}(\overline{z})$. For a geometric description we think of $X_{\text{\tiny{L}}}^{I}$ and $X_{\text{\tiny{R}}}^{I}$ as the holomorphic and antiholomorphic components of the compact bosons describing the torus, while the ${\cal X}_{\text{\tiny{L}}}^{a}$ can be thought of as representing the heterotic worldsheet (Weyl) fermions.444 The full heterotic string contains, in addition, the right-moving fermion $\psi^{I}_{\text{\tiny{R}}}(\overline{z})$—these are the superpartners of $X^{I}_{\text{\tiny{R}}}$, as well as the Minkowski degrees of freedom for ${\mathbb{R}}^{1,9-d}$ and the $bc$$\beta\gamma$ ghost system. These degrees of freedom (and the accompanying right-moving GSO projection) will not play a role in our discussion.
@@ 3 General shift orbifolds
From the preceding section, we see that the phase $U$ belongs to $\operatorname{Hom}(\Gamma,\operatorname{U{}}(1))$. Because $\Gamma$ is a self-dual lattice, any such homomorphism takes the form
@@ 3.1 Orbifold consistency conditions
Having described the most general symmetry we wish to consider, we now turn to gauging $G$, which is accomplished by introducing twisted sectors and taking a projection onto $G$–invariant states. In the heroic days the most concrete check of the consistency of such a procedure was obtained by calculating the partition function of the orbifolded theory and checking its modular invariance. In certain simple situations one could use modular orbits to construct a modular–invariant partition function and then interpret it as a sum over twisted sectors. [^5] A lucid early review of this can be found in [14]; a modern perspective is given in [1]. The more modern categorical perspective gives a more satisfying in-principle answer: we can think of the gauging as an introduction of a background field for the discrete group $G$, and there is a certain class $\omega\in H^{3}(G,\operatorname{U{}}(1))$, which must vanish if the sum over all gauge configurations is to be well-defined. [^6] A useful discussion and additional references can be found in [15], which also gives a generalization of the construction to higher dimensions.
We can follow a less technical approach because there is a particularly simple way to construct the unprojected twisted Hilbert spaces for each $\ell\in G$ based on spectral flow. At any point in the moduli space the CFT has a Kac-Moody symmetry $(\operatorname{U{}}(1)^{d+16})_{\text{\tiny{L}}}\times(\operatorname{U{}}(1)^{d})_{\text{\tiny{R}}}$, which acts on the vertex operators ${\cal V}_{{\boldsymbol{p}}}$ labeled by ${\boldsymbol{p}}\in\Gamma$. For any $\ell\in G$ we can find a linear combination of the currents, with left- and right-moving components $J^{\ell}_{\text{\tiny{L}}}$ and $J^{\ell}_{\text{\tiny{R}}}$ such that the charge of ${\cal V}_{{\boldsymbol{p}}}$ with respect to this current is precisely $\sum_{a}\ell_{a}s_{a}\cdot{\boldsymbol{p}}$. We can use these currents to construct a twist field $\Sigma_{\ell}$, such that for every vertex operator ${\cal V}_{{\boldsymbol{p}}}$ in the original Hilbert space ${\cal H}_{0}$, we have
@@
This is exactly the monodromy for the vertex operators in the twisted sector twisted by $\ell$, and we can therefore describe the twisted Hilbert space ${\cal H}_{\ell}$: it is isomorphic to ${\cal H}_{0}$, and the corresponding vertex operators are ${\cal V}_{\widetilde{{\boldsymbol{p}}}}$, where $\widetilde{{\boldsymbol{p}}}={\boldsymbol{p}}+\sum_{a}\ell_{a}s_{a}$, and ${\boldsymbol{p}}\in\Gamma$. The spin and weight of each ${\cal V}_{\widetilde{{\boldsymbol{p}}}}$ are determined by (2.10) and (2.11) with the replacement ${\boldsymbol{p}}\to\widetilde{{\boldsymbol{p}}}$, so that we know the full twisted sector spectrum. We set
@@
and we keep all operators with $\widetilde{{\boldsymbol{p}}}\in\Gamma^{\ell}$ satisfying $Q^{\ell}_{a}(\widetilde{{\boldsymbol{p}}})\in{\mathbb{Z}}$ for all $a$. We constrain these charges by the following consistency conditions.
@@
Using the distinguished basis described above, we see that
@@
This shows $v\in\Gamma_{\text{inv}}$ since each sum belongs separately to $\Gamma_{\text{inv}}$
@@ 3.3 The quantum symmetry
where ~$\widetilde{\ell}\in\widetilde{G}$. This action projects out the twisted sector states in $\Gamma^{\ell}_{\text{inv}}$ for $\ell\neq 0$ and introduces back the sectors generated by ${\boldsymbol{\sigma}}^{\ast a}$ projected out by the original orbifold.
@@ 3.4 Dualities among shift vectors
with two positive integers $k$ and $\hat{k}$. There are two possible cyclic orbifolds of the CFT using this ${\boldsymbol{\sigma}}$ allowed by the consistency conditions (3.16): a ${\mathbb{Z}}_{k}$ shift orbifold with $s={\boldsymbol{\sigma}}/k$, and a ${\mathbb{Z}}_{\hat{k}}$ shift orbifold with $\hat{s}={\boldsymbol{\sigma}}/\hat{k}$. Remarkably, the two orbifold CFTs are equivalent. The isomorphism $\varphi$ between the corresponding lattices is a reflection through the hyperplane orthogonal to ${\boldsymbol{\sigma}}$ [^7] We thank G. Chenevier for this observation, which is found in proposition 3.1.14 of [17].
@@
When the lattice $\Gamma$ is Euclidean, the above statement implies that every ${\mathbb{Z}}_{k}$ shift orbifold with shift vector $s$ admits a dual presentation as a ${\mathbb{Z}}_{\hat{k}}$ shift orbifold, where $\hat{k}=k\,s\cdot s/2$. The associated shift vector is
@@ 4.1 A nine-dimensional perspective
We now examine the $r\to 0$ limit. As addressed in [4], taking $r\to 0$ does not always provide a decompactification limit of the heterotic moduli space: it can also lead to periodic or chaotic behavior, depending on the value of the parameter ${{A}}$. In our case, coming back to (2.11), we see that any ${\boldsymbol{p}}$ satisfying $\text{n}=\frac{s\cdot s}{2}\text{w}+s\cdot{{\boldsymbol{L}}}$ will give an asymptotically holomorphic light state of spin $s({\boldsymbol{p}})=\frac{1}{2}({{\boldsymbol{L}}}+\text{w}s)\cdot({{\boldsymbol{L}}}+\text{w}s)$. Elements of the invariant lattice $\Lambda_{\text{inv}}=\cup_{\ell=0}^{k-1}\Lambda^{\ell}_{\text{inv}}$ provide us with such states. Indeed, to any $\widetilde{{{\boldsymbol{L}}}}={{\boldsymbol{L}}}+\ell s\in\Lambda^{\ell}_{\text{inv}}$ we can associate the lattice vector ${\boldsymbol{p}}=\ell{{\boldsymbol{e}}}+(s\cdot{{\boldsymbol{L}}}+\frac{\ell s\cdot s}{2}){{\boldsymbol{e}}}^{\ast}+{{\boldsymbol{L}}}\in\Gamma$ [^10] Using the description (3.21) of $\Lambda_{\text{inv}}$, we can write $\widetilde{{{\boldsymbol{L}}}}={{\boldsymbol{L}}}_{0}+\ell t$, where ${{\boldsymbol{L}}}_{0}\in\Lambda^{0}_{\text{inv}}$ and $t=s+k\zeta{\boldsymbol{\sigma}}^{\ast}$. We then see that $s\cdot{{\boldsymbol{L}}}+\ell s\cdot s/2=s\cdot{{\boldsymbol{L}}}_{0}+\ell(\zeta+s\cdot s/2)$ is an integer, due to the orbifold consistency conditions (3.15). In the $r\to 0$ limit, the corresponding vertex operator becomes holomorphic, as its right-moving weight is simply
@@ 4.2 The heterotic moduli space
where the T-duality group $\operatorname{O{}}(\Gamma)$ is identified with the isometry group of the lattice $\Gamma$ [^12] In addition to the lattice action $\varphi_{g}:\Gamma\to\Gamma$, the duality $g\in\operatorname{O{}}(\Gamma)$ must also be specified by a choice of phase $U(g,{\boldsymbol{p}})$, as we recalled in subsection 2.3. For the generators of $\operatorname{O{}}(\Gamma)$ listed below, a consistent choice of such cocyles can be found in [13]. Since they play a spectator role in our study, we do not include them here. Three subgroups of $\operatorname{O{}}(\Gamma)$ will be relevant for our analysis.
@@
    The factorized duality transformation is the involution defined as [^13] We use a slightly different terminology than in [13], by combining the factorized duality of [13] with a circle reflection. This removes an unnecessary minus sign in the transformation law of ${{A}}$.
@@
We are now ready to describe the parameter space $\mathfrak{M}_{1,17}$, following [4]. Consider the $\operatorname{Spin}(32)/{\mathbb{Z}}_{2}$ frame and decompose the Narain lattice as $\Gamma=\Lambda_{1,1}+\Lambda_{16}$. We construct the space $\mathfrak{M}_{1,17}$ by picking, inside $\operatorname{Gr}(1,17)$, a fundamental region for the action of the T-duality group $\operatorname{O{}}(\Gamma)$. This is achieved in several steps. First, using a Wilson line shift by some lattice element $\boldsymbol{q}\in\operatorname{\mathfrak{spin}}(32)\subset\Lambda_{16}$, we can bring the parameter ${{A}}$ to the maximal torus of $\operatorname{Spin}(32)$. Then we apply a Weyl symmetry that takes ${{A}}$ to a fundamental alcove of $\operatorname{Spin}(32)$—this is a fundamental domain for the action of the Weyl group on the maximal torus. Making the standard choice for this alcove, we can impose the conditions ${\boldsymbol{\beta}}_{i}\cdot{{A}}\geq 0$ and $-{\boldsymbol{\beta}}_{0}\cdot{{A}}\leq 1$ on the Weyl-transformed Wilson line parameter. The Wilson line shift by $\boldsymbol{w}_{+}$ can be combined with the Weyl transformation $R(a_{1},\dots,a_{16})=(-a_{16},\dots,-a_{1})$ to obtain a transformation that preserves the fundamental alcove of $\operatorname{Spin}(32)$. Making use of this duality, we can impose on ${{A}}$ the additional condition $\boldsymbol{w}_{+}\cdot{{A}}\leq 2$. All together, the above requirements provide us with a description of the $\operatorname{Spin}(32)/{\mathbb{Z}}_{2}$ moduli space as
@@
The heterotic moduli space can also be described in a $(\operatorname{E}_{8}\times\operatorname{E}_{8})\rtimes{\mathbb{Z}}_{2}$ frame, where the Narain lattice is $\Gamma=\Gamma_{1,1}+\Lambda_{8}+\Lambda_{8}$. In parallel with the previous discussion, we first bring ${{A}}$ in the maximal torus of $\operatorname{E}_{8}\times\operatorname{E}_{8}$ using a Wilson line shift by some $\boldsymbol{q}\in\Lambda_{8}+\Lambda_{8}$, and then to the fundamental alcove, in which ${{A}}$ has positive inner product with all the simple roots, and inner product with the highest root of each $\operatorname{E}_{8}$ factor bounded by $1$. The ${\mathbb{Z}}_{2}$ outer automorphism of $\operatorname{E}_{8}\times\operatorname{E}_{8}$ acts non-trivially on ${{A}}$, and we can use it to impose one last constraint, which we take to be $-{\boldsymbol{\alpha}}_{0}\cdot{{A}}\geq-{\boldsymbol{\alpha}}^{\prime}_{0}\cdot{{A}}$. All together, we obtain the $(\operatorname{E}_{8}\times\operatorname{E}_{8})\rtimes{\mathbb{Z}}_{2}$ moduli space
@@
Several comments are in order regarding the above description of $\mathfrak{M}_{1,17}$. First, inside the Narain moduli space the Wilson line parameter ${{A}}$ has its norm bounded by ${{A}}\cdot{{A}}\leq 2$. To see this, consider the $\operatorname{Spin}(32)/{\mathbb{Z}}_{2}$ frame. Since $\mathfrak{M}_{\operatorname{Spin}(32)/{\mathbb{Z}}_{2}}$ is a convex polytope, any point ${{A}}$ of maximal norm must be an extreme point, i.e. one that cannot be written as a convex combination of two distinct points. There are $17$ possible candidates—the vertices of the polytope. A short computation shows that only one vertex reaches the maximal norm. We denote this distinguished point as ${\mathfrak{m}}_{\text{\tiny{S}}}$. Its coordinates are given by

### v2

@@ 2.1 Lattice set up
Specializing to $d=1$, we denote the heterotic moduli by a pair (r,A)(r,{{A}}), where ${{A}}\in{\mathbb{R}}^{16}$ is the set of Wilson line parameters. A choice of (r,A)(r,{{A}}) determines a point $\rho$ in the Grassmannian $\operatorname{Gr}(1,17)$ through the orthogonal basis for ${\mathbb{R}}^{1,17}$ consisting of orthogonal vectors $\widetilde{{\pi}}$} , $\pi$}, and ${\pi}^{\circ}_{\alpha}$, $\alpha=1,\ldots,16$
@@ 2.2 Vertex operators
[…]
@@ 3 General shift orbifolds
From the preceding section, we see that the phase $U$ belongs to $\operatorname{Hom}(\Gamma,\operatorname{U{}}(1))$. Because $\Gamma$ is a self-dual lattice, any such homomorphism takes the form [^5] To see this, choose a basis $\{{{\boldsymbol{e}}}_{i}\}$ of $\Gamma$ with dual basis $\{{{\boldsymbol{e}}}^{\ast i}\}$. Any homomorphism $U\in\operatorname{Hom}(\Gamma,\operatorname{U{}}(1))$ is fixed by its values $U({{\boldsymbol{e}}}_{i})=e^{2\pi\mathrm{i}x_{i}}$ with $x_{i}\in{\mathbb{R}}$, and may be written as $U({\boldsymbol{p}})=e^{2\pi\mathrm{i}s\cdot{\boldsymbol{p}}}$ with $s=\sum_{i}x_{i}\,{{\boldsymbol{e}}}^{\ast i}$.
@@ 3.1 Orbifold consistency conditions
Having described the most general symmetry we wish to consider, we now turn to gauging $G$, which is accomplished by introducing twisted sectors and taking a projection onto GG–invariant states. In the heroic days the most concrete check of the consistency of such a procedure was obtained by calculating the partition function of the orbifolded theory and checking its modular invariance. In certain simple situations one could use modular orbits to construct a modular–invariant partition function and then interpret it as a sum over twisted sectors. [^6] A lucid early review of this can be found in [14]; a modern perspective is given in [1]. The more modern categorical perspective gives a more satisfying in-principle answer: we can think of the gauging as an introduction of a background field for the discrete group $G$, and there is a certain class $\omega\in H^{3}(G,\operatorname{U{}}(1))$, which must vanish if the sum over all gauge configurations is to be well-defined. [^7] A useful discussion and additional references can be found in [15], which also gives a generalization of the construction to higher dimensions.
We can follow a less technical approach because there is a particularly simple way to directly construct the unprojected twisted Hilbert spaces for each $\ell\in G$ based on spectral flow. Although it is not as general, this construction has the advantage over the approach based on modular orbits of the partition function: the Hilbert space interpretation of the twisted sector operators is built–in from the start, and there is no need to check positivity conditions on expansion coefficients in a putative partition function. Moreover, since we will verify that the resulting vertex operators are labeled by an even self-dual lattice, we can be sure that the operators are mutually local and have an associative OPE that closes.
At any point in the moduli space the CFT has a Kac-Moody symmetry $(\operatorname{U{}}(1)^{d+16})_{\text{\tiny{L}}}\times(\operatorname{U{}}(1)^{d})_{\text{\tiny{R}}}$, which acts on the vertex operators ${\cal V}_{{\boldsymbol{p}}}$ labeled by ${\boldsymbol{p}}\in\Gamma$. For any $\ell\in G$ we can find a linear combination of the currents, with left- and right-moving components $J^{\ell}_{\text{\tiny{L}}}$ and $J^{\ell}_{\text{\tiny{R}}}$ such that the charge of ${\cal V}_{{\boldsymbol{p}}}$ with respect to this current is precisely $\sum_{a}\ell_{a}s_{a}\cdot{\boldsymbol{p}}$. We can use these currents to construct a twist field $\Sigma_{\ell}$, such that for every vertex operator ${\cal V}_{{\boldsymbol{p}}}$ in the original Hilbert space ${\cal H}_{0}$, we have
@@
[…]
@@
[…]
@@
Because the additive properties of the sets $\Gamma^{\ell}$ are preserved by the orbifold projection, $\Gamma_{\text{inv}}$ forms a lattice, of the same rank and signature as $\Gamma$. Using the distinguished basis described above, we see that
@@
This shows $v\in\Gamma_{\text{inv}}$ since each sum belongs separately to $\Gamma_{\text{inv}}$ .
@@ 3.3 The quantum symmetry
where $\widetilde{\ell}\in\widetilde{G}$. This action projects out the twisted sector states in $\Gamma^{\ell}_{\text{inv}}$ for $\ell\neq 0$ and introduces back the sectors generated by ${\boldsymbol{\sigma}}^{\ast a}$ projected out by the original orbifold.
@@ 3.4 Dualities among shift vectors
with two positive integers $k$ and $\hat{k}$. There are two possible cyclic orbifolds of the CFT using this $\boldsymbol{\sigma}$} allowed by the consistency conditions (3.16): a ${\mathbb{Z}}_{k}$ shift orbifold with $s={\boldsymbol{\sigma}}/k$, and a $^{\mathbb{Z}}_{\hat{k}}$ shift orbifold with $\hat{s}={\boldsymbol{\sigma}}/\hat{k}$. Remarkably, the two orbifold CFTs are equivalent. The isomorphism $\varphi$ between the corresponding lattices is a reflection through the hyperplane orthogonal to $\boldsymbol{\sigma}$} [^8] We thank G. Chenevier for this observation, which is found in proposition 3.1.14 of [17].
@@
When the lattice $\Gamma$ is Euclidean, the above statement implies that every ${\mathbb{Z}}_{k}$ shift orbifold with shift vector $s$ admits a dual presentation as a $^{\mathbb{Z}}_{\hat{k}}$ shift orbifold, where $\hat{k}=k\,s\cdot s/2$. The associated shift vector is
@@ 4.1 A nine-dimensional perspective
We now examine the $r\to 0$ limit. As addressed in [4], taking $r\to 0$ does not always provide a decompactification limit of the heterotic moduli space: it can also lead to periodic or chaotic behavior, depending on the value of the parameter A{{A}}. In our case, coming back to (2.11), we see that any $\boldsymbol{p}$} satisfying $\text{n}=\frac{s\cdot s}{2}\text{w}+s\cdot{{\boldsymbol{L}}}$ will give an asymptotically holomorphic light state of spin $s({\boldsymbol{p}})=\frac{1}{2}({{\boldsymbol{L}}}+\text{w}s)\cdot({{\boldsymbol{L}}}+\text{w}s)$. Elements of the invariant lattice $\Lambda_{\text{inv}}=\cup_{\ell=0}^{k-1}\Lambda^{\ell}_{\text{inv}}$ provide us with such states. Indeed, to any $\widetilde{{{\boldsymbol{L}}}}={{\boldsymbol{L}}}+\ell s\in\Lambda^{\ell}_{\text{inv}}$ we can associate the lattice vector ${\boldsymbol{p}}=\ell{{\boldsymbol{e}}}+(s\cdot{{\boldsymbol{L}}}+\frac{\ell s\cdot s}{2}){{\boldsymbol{e}}}^{\ast}+{{\boldsymbol{L}}}\in\Gamma$. [^11] Using the description (3.21) of $\Lambda_{\text{inv}}$, we can write $\widetilde{{{\boldsymbol{L}}}}={{\boldsymbol{L}}}_{0}+\ell t$, where ${{\boldsymbol{L}}}_{0}\in\Lambda^{0}_{\text{inv}}$ and $t=s+k\zeta{\boldsymbol{\sigma}}^{\ast}$. We then see that $s\cdot{{\boldsymbol{L}}}+\ell s\cdot s/2=s\cdot{{\boldsymbol{L}}}_{0}+\ell(\zeta+s\cdot s/2)$ is an integer, due to the orbifold consistency conditions (3.15). In the $r\to 0$ limit, the corresponding vertex operator becomes holomorphic, as its right-moving weight is simply
@@ 4.2 The heterotic moduli space
where the T-duality group $\operatorname{O{}}(\Gamma)$ is identified with the isometry group of the lattice $\Gamma$. [^13] In addition to the lattice action $\varphi_{g}:\Gamma\to\Gamma$, the duality $g\in\operatorname{O{}}(\Gamma)$ must also be specified by a choice of phase $U(g,{\boldsymbol{p}})$, as we recalled in subsection 2.3. For the generators of $\operatorname{O{}}(\Gamma)$ listed below, a consistent choice of such cocyles can be found in [13]. Since they play a spectator role in our study, we do not include them here. Three subgroups of $\operatorname{O{}}(\Gamma)$ will be relevant for our analysis.
@@
    The factorized duality transformation is the involution defined as [^14] We use a slightly different terminology than in [13], by combining the factorized duality of [13] with a circle reflection. This removes an unnecessary minus sign in the transformation law of A{{A}}.
@@
We are now ready to describe the parameter space $\mathfrak{M}_{1,17}$, following [4]. Consider the $\operatorname{Spin}(32)/{\mathbb{Z}}_{2}$ frame and decompose the Narain lattice as $\Gamma=\Lambda_{1,1}+\Lambda_{16}$. We construct the space $\mathfrak{M}_{1,17}$ by picking, inside $\operatorname{Gr}(1,17)$, a fundamental region for the action of the T-duality group $\operatorname{O{}}(\Gamma)$. This is achieved in several steps. First, using a Wilson line shift by some lattice element $\boldsymbol{q}\in\operatorname{\mathfrak{spin}}(32)\subset\Lambda_{16}$, we can bring the parameter A{{A}} to the maximal torus of $\operatorname{Spin}(32)$. Then we apply a Weyl symmetry that takes A{{A}} to a fundamental alcove of $\operatorname{Spin}(32)$—this is a fundamental domain for the action of the Weyl group on the maximal torus. Making the standard choice for this alcove, we can impose the conditions ${\boldsymbol{\beta}}_{i}\cdot{{A}}\geq 0$ and ${\boldsymbol{\beta}}_{0}\cdot{{A}}\leq 1$ on the Weyl-transformed Wilson line parameter. The Wilson line shift by $\boldsymbol{w}_{+}$ can be combined with the Weyl transformation $R(a_{1},\dots,a_{16})=(-a_{16},\dots,-a_{1})$ to obtain a transformation that preserves the fundamental alcove of $\operatorname{Spin}(32)$. Making use of this duality, we can impose on A{{A}} the additional condition $\boldsymbol{w}_{+}\cdot{{A}}\leq 2$. All together, the above requirements provide us with a description of the $\operatorname{Spin}(32)/{\mathbb{Z}}_{2}$ moduli space as
@@
The heterotic moduli space can also be described in a $(\operatorname{E}_{8}\times\operatorname{E}_{8})\rtimes{\mathbb{Z}}_{2}$ frame, where the Narain lattice is $\Gamma=\Gamma_{1,1}+\Lambda_{8}+\Lambda_{8}$. In parallel with the previous discussion, we first bring A{{A}} in the maximal torus of $\operatorname{E}_{8}\times\operatorname{E}_{8}$ using a Wilson line shift by some $\boldsymbol{q}\in\Lambda_{8}+\Lambda_{8}$, and then to the fundamental alcove, in which A{{A}} has positive inner product with all the simple roots, and inner product with the highest root of each $\operatorname{E}_{8}$ factor bounded by $1$. The ${\mathbb{Z}}_{2}$ outer automorphism of $\operatorname{E}_{8}\times\operatorname{E}_{8}$ acts non-trivially on A{{A}}, and we can use it to impose one last constraint, which we take to be ${\boldsymbol{\alpha}}_{0}\cdot{{A}}\geq-{\boldsymbol{\alpha}}^{\prime}_{0}\cdot{{A}}$. All together, we obtain the $(\operatorname{E}_{8}\times\operatorname{E}_{8})\rtimes{\mathbb{Z}}_{2}$ moduli space
@@
Several comments are in order regarding the above description of $\mathfrak{M}_{1,17}$. First, inside the Narain moduli space the Wilson line parameter A{{A}} has its norm bounded by ${{A}}\cdot{{A}}\leq 2$. To see this, consider the $\operatorname{Spin}(32)/{\mathbb{Z}}_{2}$ frame. Since $\mathfrak{M}_{\operatorname{Spin}(32)/{\mathbb{Z}}_{2}}$ is a convex polytope, any point A{{A}} of maximal norm must be an extreme point, i.e. one that cannot be written as a convex combination of two distinct points. There are $17$ possible candidates—the vertices of the polytope. A short computation shows that only one vertex reaches the maximal norm. We denote this distinguished point as ${\mathfrak{m}}_{\text{\tiny{S}}}$. Its coordinates are given by